OpenStudy (anonymous):

The number of computers sold by BCC depends on the dollar amount, x, that they spend on advertising. How much must they spend to sell 300 computers? Round to the nearest dollar and do not include units in your answer.

OpenStudy (anonymous):

N(x) = 100 + 20 * ln(0.25x) thats the equation your supposed to use, i think

OpenStudy (anonymous):

300 = 100 + 20ln(x/4) 10 = ln(x/4) x/4 = e^10

OpenStudy (anonymous):

so is it 88105.86?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

YES!!!!!!! it was right. thank you thank you